Let G={3^n;n element of Z} prove that (G,•)
is an abelian group
1
Expert's answer
2020-09-24T18:07:51-0400

Given "G=\{3^n : n \ element \ of \ \Z\}" .

Closure: Let any "a\in G, b\in G" , so "a = 3^n, b = 3^m" where "n \in \Z, m\in \Z" .

Now "a \times b = 3^n \times 3^m = 3^{n m}" where "n m \in \Z" "\implies ab \in G" .

Associative: Let "a, b, c \in G" , so "a = 3^n, b = 3^m, c = 3^k" ,

hence "(a \times b) \times c = (3^{nm}) \times 3^k = 3^{nmk} = 3^n \times (3^m \times 3^k) = a \times (b\times c)" .

Identity: For all "a \in G" , "a = 3^m" where "m \in \Z" .

There exists "e = 1 = 3^0" where "0 \in \Z" , such that "a \times e = e\times a = a" .

Inverse: For all "a \in G" , "a = 3^m" where "m \in \Z" .

There exists "b = 3^{-m} : a\times b = b\times a = 1" .

Commutative: "a \times b = 3^{nm} = b \times a \ \ \forall \ a, b \in G" .